Alfredo has 3 times as many 5-peso coins as Antonio. If Alfredo gives Antonio 6 coins, Alfredo will be left with 4 more coins than Antonio. What is the total number of 5-peso coins that Alfredo and Antonio have?
Answer by richwmiller(17219)

r=3*n
r-6=n+4+6
r=n+16
3n=n+16
2n=16
n=8 Antonio
r=24 Alfredo
24-6=8+6
18=14+4